5. Joe Jones, Inc. has a beta of .85. The risk-free rate is 5% and the expected rate of return on the market portfolio is 10%. a. Compute the required return for Joe Jones using the security market line (SML) equation. b. What is “beta?” Under what rationale is beta an appropriate measure of risk?
Answer #1
Calculation of required return:
Required return= Risk free rate+ beta*(market return- risk free rate)
Required return= 5+0.85*(10-5)
                                = 5+4.25=9.25%
Required return= 9.25%